when a dielectric slab of thickness 6 cm is introduced between the plates of parallel plate condenser, it is found that the distance between the plates has to be increased by 4 cm to restore to capacity to original value . the dielectric constant of the slab is ?

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
let Co=eA/D
let c1= keA/(D+6)

keA/(D+6+10)=eA/(D)
k=(D+16)/D

sher mohammad
askiitians faculty
b.tech, iit delhi